What is the pH after 0.29 moles of NaOH is added to a 1.00L buffer containing 1.00 moles of NH3 and 1.00 moles NH4Cl ?

Kb NH3 = 1.8 x 10-5
Answer #1

moles of Naon = 0.29 mol volume (total) =1h Molarity of Naon = moles of Naon volume in h 0.29 mol -0.29M 1 h Molarity of NH3. por = -log, 8 + 5 logio + logo. 71 - log1:29 = -0.255 3 + 5.+ 60.1487) - 0.1106 pon = 4.4854 pH= = 14-4.4854 9.5146 [pH = 9